Using the data below, answering the following questions:                                        

Fixed costs         $12,500,000                       

Variable cost/day            $350                      

Charge (rev)/day             $1,200                  

Inpatient days   16,000                  

  1. Construct the hospital's base case projected P&L statement                                      
  2. What is the hospital's breakeven point (volume / patient days needed to breakeven)?            
  3. What is the economic breakeven (volume required) to provide a profit of $1,000,000?  
  4. What is the total contribution margin if volume decreases by 20%?

Based on the scenario in d., if fixed costs remain the same, what is the hospital's profit or loss?
Answer #1

A.Construct the hospital's base case projected P&L statement :

Revenue - Variable Cost = Contribution .

Contribution- Fixed Cost = Profit

Total Revenue 16000*1200 19200000
Total Variable Cost 16000*350 5600000
Contribution Margin 16000*(1200-350) 13600000
Fixed Cost 12500000
Profit 1100000

B.What is the hospital's breakeven point (volume / patient days needed to breakeven)?   

Break Even Point = Fixed costs / ( selling price – variable costs )

12500000/(1200-350) = 14705.88 Inpatient Days.

C.What is the economic breakeven (volume required) to provide a profit of $1,000,000?  

The formula for calculation of output to earn a certain amount of profit is as follows :

(Fixed Cost + Desired Profit) / Contribution Per Unit

(12500000+1000000) / (1200-350) = 15882.35 Inpatient Days.

D.What is the total contribution margin if volume decreases by 20%?

Here Inpatient Days is Volume . i.e. 16000 . Volume decreases by 20% i. e 3200

Volume = 16000-3200 = 12800 Inpatient Days

Total Revenue 12800*1200 15360000
Total Variable Cost 12800*350 4480000
Contribution Margin 12800*(1200-350) 10880000

Based on the scenario in d., if fixed costs remain the same, what is the hospital's profit or loss?

Profit / Loss = Contribution Margin - Fixed Cost

10880000-12500000 = -1620000.

i.e. Loss is 1620000.